OpenStudy (anonymous):

Every day, there are 2 times more likes on an internet video of a bird which is modeled by the function c(n) = (2)^(n − 1), where n is the number of days since the video posted. On the first day, there were 122 likes. What is the function that shows the number of likes each day?

OpenStudy (anonymous):

c(n) = (2)^(122 − 1) c(n) = 122(2)^(n − 1) c(n) = (2)^(122)(n − 1) c(n) = (122)^(n − 1)

OpenStudy (anonymous):

This site helps me. Might wanna give it a try it breaks down the problem http://www.wolframalpha.com/input/?i=%E2%88%92x2+%2B+5x+%2B+6+%3D+3x+%E2%88%92+2%3F

OpenStudy (anonymous):

I don't know how that would help me here @viacomplys

OpenStudy (turingtest):

check each one c(n) = (2)^(122 − 1) according to the formula, how many likes will this have on day 1?

OpenStudy (anonymous):

That one would have 2.7 * 10^36

OpenStudy (anonymous):

So obviously wrong, considering that is a really really big number

OpenStudy (anonymous):

Well, starts with 122 likes after the first day so naturally for n=1 , c(n)=122. Then it says the number of likes doubles every day so for n=2 we have 244 likes, n=3 we have 488 likes and so forth which indicates an exponential growth. So the answer, outside the context of those notations I am not familiar with, should be something like c(n)=(2^(n-1))*122 in order to verify the context.

OpenStudy (turingtest):

correct, so the next one would be c(n) = 122(2)^(n − 1) how many people does it have on day 1?

OpenStudy (anonymous):

122, since 2^0 is 1 and 122*1 is obviously 122.

OpenStudy (turingtest):

right what about day 2?

OpenStudy (anonymous):

would be 244, right?

OpenStudy (turingtest):

yep, and for day 3?

OpenStudy (anonymous):

488?

OpenStudy (anonymous):

Well, starts with 122 likes after the first day so naturally for n=1 , c(n)=122. Then it says the number of likes doubles every day so for n=2 we have 244 likes, n=3 we have 488 likes and so forth which indicates an exponential growth. So the answer, outside the context of those notations I am not familiar with, should be something like c(n)=(2^(n-1))*122 in order to verify the context.

OpenStudy (turingtest):

correct, so what do you conclude?

OpenStudy (anonymous):

the answer is B

OpenStudy (turingtest):

I agree :)

OpenStudy (anonymous):

Thank you so much :)